An object with a velocity of 1m/s accelerates with an acceleration of 1m/s/s. What is its velocity after moving 1m ?P. S: the answer is 1.7m/s

Grade:12

1 Answers

Rishav Ranjan
6 years ago
Use the third equation of motion=v^2=u^2+2asu=1m/s and a=1m/s^2Substituting the value we have;v^2=1+2×1×1=3v=√3=1.7(approx)